OpenStudy (angelinaxoxo1):

Write a subtraction equation, and a multiplication equation that has -2 as a solution. I"LL FAN AND MEDAL

OpenStudy (jtug6):

Simple. Subtraction: a-b = -2 Multiplication: x*y=-2 Find values for a/b/x/y, unless they require you to use the same variables for both equations.

OpenStudy (angelinaxoxo1):

How is that any help?

OpenStudy (jtug6):

Because you need two equations. One has to be subtraction (top one) and multiplication (bottom one). You can literally pick any numbers you want for the variables (a,b,x,y) so long as the equation when solved equals -2. For example, 0 - 2 = -2. So, use a = 0, and b = 2 perhaps? Is that clearer?

OpenStudy (angelinaxoxo1):

Sure.

OpenStudy (jtug6):

And same for multiplication. What times what can give you -2? Well, anything times 1 is still the same. So, maybe x can be -2 and y can be 1. =)

OpenStudy (robtobey2):

\[\text{Solve}[\{x+y=-2,x* y=-2\}] \]\[\left\{x=-1-\sqrt{3},y=\sqrt{3}-1\right\},\left\{x=\sqrt{3}-1,y=-1-\sqrt{3}\right\} \]